What is the equation of the line that passes through(0-4) with a slope of 8?
Answer by venugopalramana(3286)

What is the equation of the line that passes through P=(X1,Y1)=(0,-4) with a slope of M=8?
Y-Y1=M(X-X1)
Y-(-4)=8(X-0)
Y=8X-4